How does the elapsed time for a process seem to be longer, an observer moving with the

process or observer moving relative to the process? Which observer measures proper

time? Explain your answer


1
Expert's answer
2021-03-01T11:51:23-0500

Explanation


  • It has found that time is seen to be shorter at relativistic speeds: speeds very near the speed of light.
  • This is distinguishable only at relativistic speeds hence in day-to-day activities this is not experienced.
  • This relationship is expressed in the equation,

"\qquad\qquad
\begin{aligned}
\small T&= \small \gamma\cdot \tau=\frac{1}{\sqrt {1-\frac{v^2}{c^2}}}\cdot\tau\\
\end{aligned}"

  • T is the time measured by the observer moving relative to the process.
  • "\small \tau" is that measured by the observer moving with the process.
  • "\gamma" is called the Lorentz factor which values larger numbers when the speed of the moving frame gets near the speed of light.


  • Proper time is measured by the observer moving with the process whereas the observer moving relative to that measures some extended time.